OpenStudy (anonymous):

Solve for x: |x| − 8 = −5 Can someone explain the steps for how to solve this please?

OpenStudy (anonymous):

the answer is 3, but if you subtract 8-5 that equals 3 which will give you the answer

OpenStudy (anonymous):

its x - 8 = -5 the five is negative tho so..

OpenStudy (anonymous):

I just want to know the steps to figuring this out cuz i know thats not the right answer sorry..

OpenStudy (mathstudent55):

First, add 8 to both sides. Can you do that?

OpenStudy (lynfran):

\[|x|-8=-5\]ok 1st step it to add 8 to both sides of the equation

OpenStudy (mathstudent55):

\(\large |x| - 8 = -~5\) \(\large ~~~~~+8~~~~~+8\)

OpenStudy (anonymous):

so 16 and 3?

OpenStudy (mathstudent55):

No. -8 + 8 = 0 -5 + 8 = 3. That is good.

OpenStudy (mathstudent55):

You end up with \(\large |x| = 3\) Ok so far?

OpenStudy (anonymous):

the 8 is not negative tho its x minus 8 equals negative 5

OpenStudy (lynfran):

example solve for x... \[|x|-7=9\]\[|x|=9+7\]\[|x|=16\]\[x=\pm 16\]

OpenStudy (mathstudent55):

In our original equation, -8 (minus eight) means subtract 8. A subtraction can always be written as the sum of the opposite. So -8 means plus the opposite of 8. \(\large |x| - 8 = -5\) \(\large |x| + (-~8) = -5\) \(\large~~~~~~~~~~+8~~~~~+8\) After adding 8 to both sides, you get: \(\large |x| = 3\)

OpenStudy (mathstudent55):

Do you understand now how we add 8 to both sides to isolate \(|x|\)?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

the options for it tho aren't just one number its one of these x = −13 and x = −3 x = 3 and x = −3 x = 3 and x = 13 No solution

OpenStudy (anonymous):

plus i dont understand how you turned minus eight into negative eight, i dont understand why you would do that cri

OpenStudy (mathstudent55):

Ok. Now we have a simple absolute value equation. On the left side we have just \(|x|\). This equals a number, 3, on the right side. Whenever you have an absolute value equation where something in absolute value signs equals a number, you separate it into two simple equations, both without absolute signs. This is the pattern. Then I'll show you with your equation. To solve the absolute value equation \(\large |x| = k\), where k is a non-zero number, solve these two equations \(\large x = k\) or \(\large x = -k\)

OpenStudy (lynfran):

heres another example ...